OpenStudy (anonymous):

1. Tom and his best friend Edward are seeking to join a gym. Tom saw on television that Platinum Gym was having a grand opening in a few months. Tom is considering a member (c) If Tom decides to join Platinum Gym and Edward joins Super Fit, who would pay less during the first month of membership? Who would pay less after a year of membership? Show your work. (d) Is there a membership period for which Tom and Edward would pay the same amount? If so, how long would this membership be and how much would it cost? Show your work. (e) Which gym membership is a better deal? Explain why. Answer:

OpenStudy (anonymous):

OK Well first is there prices of the gyms that we can compare?

OpenStudy (anonymous):

Well theres only this 1. om and his best friTend Edward are seeking to join a gym. Tom saw on television that Platinum Gym was having a grand opening in a few months. Tom is considering a membership at Platinum Gym. Platinum Gym charges a one-time registration fee of $60, and membership is $20 a month. Tom’s best friend Edward says he plans to join Super Fit because Super Fit only charges $10 a month but has a one-time registration fee of $150.

OpenStudy (anonymous):

*friend

OpenStudy (anonymous):

Sorry so sorry guys i got the answer!

OpenStudy (anonymous):

OK good job

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Your Welcome! If you do need help with anything else just ask!

OpenStudy (anonymous):

60 + 20(1) = 60 + 20 = 80 per month for Tom at Platinum gym 150 + 10(1) = 150 + 10 = 160 per month for Edward at Super fit so for 1 month, Tom pays less for a year.. Tom pays : 80 * 12 = 960 Edward pays : 160 * 12 = 1920 Tom still pays less...even at 1 yr ======================== 60 + 20m = 150 + 10m 20m - 10m = 150 - 60 10m = 90 m = 9 --- they would pay the same at 9 months 60 + 20(9) = 60 + 180 = 240 is what they would both pay ========================= The better deal depends on how long they want the membership. If the membership is LESS then 9 months, then Tom has the better deal at Platinum..but if the membership is MORE then 9 months, then Edward has the better deal at Superfit.